a plane is flying against the wind and it takes 6 hours then it flies back with the wind and it takes 4 hours. the plane is going 90 miles and hour. find the spped of the wind
Answer by mananth(16946)

speed of plane in still air = 90 mph
let wind speed = x mph
..
against wind speed = 90-x
with wind speed = 90+x
..
time against wind = 6 hours
time with wind = 4 hours
..
d= rt
..
6(90-x)=4(90+x)
540-6x=360+4x
6x+4x=540-360
10x=180
/10
x= 18 mph the speed of wind.
...
m.ananth@hotmail.ca